# Semiclassical Backreaction: A Qualitative Assessment

### v1

@@ Appendix B Numerics and performance tests
To find a numerical approximation of this solution we need to discretize space on a grid, writing $x_{i}=i\Delta x$ and $y_{i}=i\Delta x$, and solve the equation in a finite domain defined to be a square box of size $L$ (centered around the origin). The number $N^{2}$ of gridpoints we include determines the spatial resolution $\Delta x=L/N$. Finally, we also discretize time, by defining $t_{i}=i\Delta t$, and evolve the wavefunction by rewriting (46) as
$\psi(t_{f},x_{f},y_{f})=\Pi_{j=0}^{f}e^{-i\frac{\Delta t}{2}H_{k}}e^{-i\Delta tH_{p}}e^{-i\frac{\Delta t}{2}H_{k}}\psi_{0}(x_{f},y_{f}),$ (47)
@@
Figure 9: The degree of conservation of the expectation value of $H$ for the previous three choices of physical parameters. In all cases energy is conserved up to 1 part in $10^{4}$.

### v2

@@ Appendix C Numerics and performance tests
To find a numerical approximation of this solution we need to discretize space on a grid, writing $x_{j}=j\Delta x$ and $y_{j}=j\Delta x$, and solve the equation in a finite domain defined to be a square box of size $L$ (centered around the origin). The number $N^{2}$ of gridpoints we include determines the spatial resolution $\Delta x=L/N$. Finally, we also discretize time, by defining $t_{j}=j\Delta t$, and evolve the wavefunction by rewriting (49) as
$\psi(t_{f},x_{j},y_{j})=\Pi_{s=0}^{f}e^{-i\frac{\Delta t}{2}H_{k}}e^{-i\Delta tH_{p}}e^{-i\frac{\Delta t}{2}H_{k}}\psi_{0}(t_{0},x_{j},y_{j}),$ (50)
@@
Figure 9: The degree of conservation of the expectation value of $H$ for the benchmark simulations of the previous three choices of physical parameters. In all cases energy is conserved up to 1 part in $10^{4}$.